solve the equations by factoring.
20x to the 2nd power + 33x + 10 = 0

Answer by checkley77(12844)

20x^2+33x+10=0
(5x+2)(4x+5)=0
5x+2=0
5x=-2
x=-2/5 answer.
4x+5=0
4x=-5
x=-5/4 anwer.